# SN2023syz and SN2025cbj: Two Type IIn Supernovae Associated with IceCube High-energy Neutrinos

### v1

@@ Abstract
Type IIn supernovae (SNe IIn) are a subclass of core-collapse SNe in which strong interactions occur between the ejecta and dense circumstellar material, creating ideal conditions for the production of high-energy neutrinos. This makes them promising candidate sources of neutrinos. In this work, we conduct an association study between 163 SNe IIn observed by the Zwicky Transient Facility and 138 neutrino alert events detected by the IceCube neutrino observatory. After excluding alerts with poor localization, we find two SNe that are spatiotemporally coincident with neutrino events. IC231027A and IC250421A coincide with the positions of SN2023syz and SN2025cbj, respectively, within their localization uncertainties, and the neutrino arrival times are delayed by 38 days and 61 days relative to the discovery times of the corresponding SNe. Using Monte Carlo simulations, we estimate that the probability of such two coincidences occurring by chance in our sample is $p\sim 0.67\%$, suggesting a high likelihood that they arise from genuine associations, though the result is not yet statistical significant. Furthermore, model calculations show that the expected numbers of neutrino events from these SNe IIn could be consistent with the actual observations. Our study provides possible evidence that interacting SNe may be potential neutrino-emitting sources.
Neutrino astronomy
@@ 1 Introduction
In addition to serving as sources of MeV neutrino emission (as has been shown by Supernova 1987A (K. Hirata et al., 1987)), supernovae have also been extensively studied as potential sources of high-energy (TeV–PeV) neutrinos (V. N. Zirakashvili & V. S. Ptuskin, 2016; K. Murase, 2018; H.-N. He et al., 2018; Z. Li, 2019; P.-W. Chang et al., 2024; R. Abbasi et al., 2023b; S. P. Cosentino et al., 2025). Among all types of supernovae, Type IIn SNe are a subclass of core-collapse supernovae (CCSNe) whose spectra exhibit strong, narrow hydrogen emission lines (“n” standing for “narrow”). These narrow lines, often accompanied by long-lasting and luminous light curves, originate from the collision of high-velocity supernova ejecta with very dense circumstellar material (CSM) shed by the progenitor star shortly before explosion (N. Smith, 2017). Such intense interaction between the supernova ejecta and the dense CSM creates ideal conditions for high-energy neutrino production. As the fast-moving supernova ejecta plows into the dense CSM, powerful shocks can form, accelerating protons to ultra-high energies (K. Murase et al., 2011). These high-energy protons then collide with the dense CSM through proton–proton ($pp$) interactions, producing muons which in turn decay into neutrinos (K. Murase et al., 2011; K. Murase, 2018; A. Kheirandish & K. Murase, 2023; T. Pitik et al., 2023). Because having both an powerful particle accelerator and a dense target, Type IIn supernovae possess the two key ingredients needed for neutrino production, making them stand out among other types of supernovae as promising sources detectable by the IceCube neutrino observatory.
@@
Models show that nearby SNe IIn (tens of Mpc or closer) with efficient CR acceleration could produce a neutrino signal detectable by IceCube (M. Petropoulou et al., 2017). As a population they might contribute a non-negligible fraction of IceCube’s diffuse flux under optimistic assumptions (M. Petropoulou et al., 2017). Single-source and stacking searches conducted with IceCube have yielded meaningful constraints, but so far there is no definitive and widely accepted detection of a Type IIn supernova as a high-energy neutrino source (R. Abbasi et al., 2023b). However, follow-up observations of neutrino alerts do find a few interacting SNe candidates (interacting SNe in general including Ibn/IIn) (M. G. Aartsen et al., 2015b; R. Stein et al., 2025). Recently, R. Stein et al. (2025) found an interacting supernova, SN 2023uqf, in the optical follow-up observations of IceCube alert events with the Zwicky Transient Facility (ZTF), which coincided in time with the high-energy neutrino IC231004A, providing a observational evidence that interacting supernovae can serve as hadronic cosmic-ray accelerators.
@@
In view of the above arguments, a systematic search for associations between SNe IIn and neutrinos is necessary. This will help determine whether SNe IIn are effective high-energy neutrino emitters and quantify their contribution to the IceCube diffuse neutrino flux. In this work, we conduct a systematic search for spatial-temporal coincidences between SNe IIn and IceCube alert events. We use the SNe IIn classified by ZTF - Bright Transient Survey (BTS) (D. A. Perley et al., 2020) as our supernova sample. This is a public catalog with a strict pipeline to conduct a classification for transient sources, and it conducts a magnitude-limited ($m<19\,{\rm mag}$ in either the $g$ or $r$ filter) survey for extragalactic transients in the ZTF public stream (D. A. Perley et al., 2020). We find two spatial-temporal coincidence events, namely IC231027A-SN2023syz ($z$ = 0.037) and IC250421A-SN2025cbj ($z$ = 0.0675). We calculate the model-predicted number of neutrino events for these two sources to check whether they are compatible with observations.
@@ 2.1 The Neutrino Sample
We note that R. Abbasi et al. (2023c) has provided alert events from before 2019. However, considering the start time of the ZTF operation, we focus primarily on neutrinos released after 2019, since this period better matches the survey time of the ZTF-BTS catalog. This allows alert events to occur after the SNe listed in the ZTF-BTS catalog. Therefore, for consistency, in this analysis we use only alerts listed in the GCN after 2019. We include all bronze and gold alert events in our sample, which have at least a 30% probability of astrophysical origin. In addition, in the association analysis, good spatial localization of neutrino events is essential, as large localization errors can lead to a high rate of chance coincidences. Many previous studies (e.g., A. Plavin et al. (2020); T. Hovatta et al. (2021)) excluded alert events when the area of their 90% confidence level (C.L.) error region (denoted as $\Omega_{90}$) exceeds a certain threshold. Accordingly, in this analysis, we retain only alert events with $\Omega_{90}\leq 30\,\text{deg}^{2}$ to remove those alerts with large error regions. Our final alert sample contains a total of 138 alert events from June 19, 2019 to June 30, 2025, as shown in Fig. 1.
@@ 2.3 Crossmatch and Chance Probability
We perform a systematic search for SNe IIn coincident with neutrino events based on the following criteria: (1) the position of the SN IIn lies within the $r_{90}$ error radius of the neutrino event; (2) the arrival time of the neutrino alert event is within 100 days of the discovery date of the SN IIn. We find that SN 2025cbj and SN 2023syz are spatially and temporally associated with the neutrino events IC250421A and IC231027A, respectively (see Table 1 for their information). IC250421A and IC231027A arrived 61 and 38 days after the discovery date of SN 2025cbj and SN 2023syz (2025 February 19 and 2023 September 19, MJD == 60725 and 60206), respectively. The angular distances between the SNe and the central positions of their corresponding alert events are 1.79 for SN 2025cbj and 2.10 for SN 2023syz.
@@
$t_{\rm rise,obs}$ Time delay $M_{\rm abs}$ Redshift RA DEC Association $E_{\nu}$
[days] [days] [deg] [deg] [TeV]
SN2023syz 10 38 17.58-17.58 0.037 268.85 45.22 IC231027A 191.5
SN2025cnj 50 61 19.15-19.15 0.0675 239.92 27.11 IC250421A 151.4
We then estimate the chance probability of obtaining two coincidence events using Monte Carlo (MC) simulations. Since IceCube is located at the South Pole and its detection sensitivity depends only on the zenith angle (M. G. Aartsen et al., 2017b, 2020), we generated mock neutrino samples by randomizing the right ascension (RA) while keeping the declination (DEC) and $r_{90}$ error radius fixed at their observed values. The arrival times of the mock neutrinos are uniformly drawn between the earliest and latest arrival times in the dataset. In addition, we generate simulated SN IIn samples by randomizing their positions within a radius of $10^{\circ}$ around their original locations, while requiring that each simulated source remain within the sky coverage of the ZTF-BTS field. We also apply a Kolmogorov–Smirnov (KS) test to the distributions of $\rm RA$ and $\rm DEC$ for the simulated sources to ensure that the mock samples preserve the distribution characteristics of the real sources.
By performing $10^{4}$ Monte Carlo simulations, we estimate the chance probability (i.e., the $p$-value for rejecting the null hypothesis) of finding at least two matches with a time delay of less than 61 days. The probability was computed using p=(M+1)/(N+1)p={(M+1)}/{(N+1)} where $M$ is the number of simulations in which the statistic (i.e., the number of matches) is greater than or equal to the observed value, and $N=10^{4}$ is the total number of MC simulations (A. C. Davison & D. V. Hinkley, 1997). We obtain a chance probability of $p\sim 0.67\%$. This result reflects that the probability of occurring two (or more) coincidence in our sample by chance is low. It is rather likely that the SNe and neutrino associations we have found are physically real, but the current results have not yet reached a level of statistically significant.
3 Model-predicted neutrino numbers from SN 2025cbj and SN 2023syz
@@
where $R_{\rm dec}={M_{\rm ej}u_{w}}/{\dot{M}}$ is the deceleration radius, and $s=10$ is the outer slope of the ejecta density profile. Consequently, the shock velocity is $v_{\rm{sh}}(t)=dR_{\rm sh}/dt\propto t^{(-1)/(s-2)}$ for $R\leq R_{\rm dec}$ and $v_{\rm{sh}}(t)\propto t^{-1/3}$ for $R>R_{\rm dec}$.
@@
where $\alpha$ is the radial dependence index of the shock velocity, $\alpha=-1/7$ (1/2-1/2) for $R<R_{\rm dec}$ ($R>R_{\rm dec}$). $\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$ are the minimum Lorentz factor of protons, the proportion of kinetic energy converted into proton acceleration, and the proton spectral index, respectively. We follow the previous works of, e.g., T. Pitik et al. (2022); K. Murase et al. (2011), and set them as ($\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$) = (1,0.1,21,0.1,-2). The maximum Lorentz factor of protons $\gamma_{p,\rm max}$ can be calculated by requiring that the acceleration timescale is shorter than the cooling timescale. For proton cooling, we consider the cooling due to inelastic collisions and adiabatic expansion of the shocked shell. More details on the calculation of $\gamma_{p,\rm max}$ can be seen in T. Pitik et al. (2022).
@@
where Aeff($E_{\nu}$, $\delta$) is the detector effective area for which we use the effective area given in Ref. (R. Abbasi et al., 2023c). The minimum neutrino energy $E_{\nu,\rm min}$ depends on the effective area, and $E_{\nu,\rm max}$ is related to the maximum protons energy. In Figure 2, we present the neutrino event rate of SN2023syz and SN2025cbj for chosen sets of model parameters. By integrating the temporal interval from the breakout time $t_{\rm{bo}}$ to the arrival dates of alert events $t_{f}$, i.e. $t_{\rm bo}=60206$ (MJD), $t_{f}=60244$ (MJD) for SN2023syz and $t_{\rm bo}=60725$ (MJD), $t_{f}=60786$ (MJD) for SN2025cbj, the total number of neutrinos can be calculated by $N_{\nu_{\mu}+\bar{\nu}_{\mu}}=\int_{t_{\rm{bo}}}^{t_{f}}dt\ \dot{N}_{\nu_{\mu}+\bar{\nu}_{\mu}}(t)$.
@@
Figure 2: Event rates for neutrinos from SN 2023syz (left panel) and SN 2025cbj (right panel) as a function of the time after the shock breakout assuming $E_{k}=10^{52}$ erg and $E_{k}=3\times 10^{52}$ erg, respectively. The blue band denotes the variation arising from varying the $M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$ parameters in the range of ($M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$) $\in$ ($1-20]\,M_{\odot}$, $1-30]\,M_{\odot}$, $1-4]\times 10^{16}\,\rm cm$) for SN2023syz and ($M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$) $\in$ ($_{\odot}, [170]M[1-70]\,M_{\odot}$, $1-4]\times 10^{16}\,\rm cm$) for SN2025cbj. When deriving the blue band, we have excluded the parameters not satisfying $t_{\rm rise}\leq t_{\rm rise,obs}\leq 1.5\times t_{\rm rise}$ and $t_{\rm CSM}-t_{\rm bo}$ greater than the observed time delay, please see the main text and Figure 3 for details.
@@
Figure 3 presents the model-predicted total number of muon (anti)neutrino events detected by IceCube from SN2023syz and SN2025cbj. We choose specific $E_{k}$ and $R_{\rm CSM}$ values and scan the parameter space of ($M_{\rm ej}$, $M_{\rm CSM}$). In the figure, we have excluded those parameters (shaded region) in the parameter space according to the $t_{\rm{rise}}$ (the time interval between the first detection time [^4] Note that using the first detection time is a conservative choice, as emission is usually present before it. and the peak time of the SNe) and $t_{\rm{CSM}}-t_{\rm{bo}}$ (the time delay between $t_{\rm{bo}}$ and $t_{\rm{CSM}}$). According to the light curves of SN2023syz [^5] https://lasair-ztf.lsst.ac.uk/objects/ZTF23abfglcy/ and SN2025cbj [^6] https://lasair-ztf.lsst.ac.uk/objects/ZTF25aagbrpb/, we adopt the $t_{\rm{rise,obs}}$ of 10 days and 50 days for SN2023syz and SN2025cbj, respectively. The $t_{\rm{rise}}$ can be linked to the other quantities through the following relation (S. Ginzburg & S. Balberg, 2012; T. Pitik et al., 2023),

### v3

@@ Abstract
Type IIn supernovae (SNe IIn) are a subclass of core-collapse SNe in which strong interactions occur between the ejecta and dense circumstellar material, creating ideal conditions for the production of high-energy neutrinos. This makes them promising candidate sources of neutrinos. In this work, we conduct an association study between 163 SNe IIn observed by the Zwicky Transient Facility and 138 neutrino alert events detected by the IceCube neutrino observatory. After excluding alerts with poor localization, we find two SNe that are spatiotemporally coincident with neutrino events. IC231027A and IC250421A coincide with the positions of SN2023syz and SN2025cbj, respectively, within their localization uncertainties, and the neutrino arrival times are delayed by 38 days and 61 days relative to the discovery times of the corresponding SNe. Using Monte Carlo simulations, we estimate that the probability of such two coincidences occurring by chance in our sample is $p\sim 0.67\%$, suggesting that they may originate from genuine physical associations, though the result is not yet statistically significant. Our model calculations, however, indicate that the likelihood of a neutrino originating from IC231027A is low, implying that the association between IC231027A and SN2023syz is likely coincidental. Nevertheless, under optimistic parameters, the probability of detecting a neutrino from the whole SNe IIn sample could reach $\gtrsim 6\%$, indicating that detecting neutrino emission from the SNe population may be possible. Our study provides a systematic analysis, combining statistical analysis and model calculations, to assess whether interacting supernovae can serve as potential sources of neutrino emission.
Neutrino astronomy (1100) — Supernova neutrinos (1666) — High energy astrophysics (739)
@@ 1 Introduction
In addition to serving as sources of MeV neutrino emission (as has been shown by Supernova 1987A (K. Hirata et al., 1987)), supernovae have also been extensively studied as potential sources of TeV–PeV neutrinos (V. N. Zirakashvili & V. S. Ptuskin, 2016; K. Murase, 2018; H.-N. He et al., 2018; Z. Li, 2019; P.-W. Chang et al., 2024; R. Abbasi et al., 2023b; S. P. Cosentino et al., 2025). Among all types of supernovae, Type IIn SNe are a subclass of core-collapse supernovae (CCSNe) whose spectra exhibit strong, narrow hydrogen emission lines (“n” standing for “narrow”). These narrow lines, often accompanied by long-lasting and luminous light curves, originate from the collision of high-velocity supernova ejecta with very dense circumstellar material (CSM) shed by the progenitor star shortly before explosion (N. Smith, 2017). Such intense interaction between the supernova ejecta and the dense CSM creates ideal conditions for high-energy neutrino production. As the fast-moving supernova ejecta plows into the dense CSM, powerful shocks can form, accelerating protons to ultra-high energies (K. Murase et al., 2011). These high-energy protons then collide with the dense CSM through proton–proton ($pp$) interactions, producing muons which in turn decay into neutrinos (K. Murase et al., 2011; K. Murase, 2018; A. Kheirandish & K. Murase, 2023; T. Pitik et al., 2023). Because having both an powerful particle accelerator and a dense target, Type IIn supernovae possess the two key ingredients needed for neutrino production, making them stand out among other types of supernovae as promising sources detectable by the IceCube neutrino observatory.
@@
Models show that nearby SNe IIn (tens of Mpc or closer) with efficient CR acceleration could produce a neutrino signal detectable by IceCube (M. Petropoulou et al., 2017). As a population they might contribute a non-negligible fraction of IceCube’s diffuse flux under optimistic assumptions (M. Petropoulou et al., 2017). Single-source and stacking searches conducted with IceCube have yielded meaningful constraints, but so far there is no definitive and widely accepted detection of a Type IIn supernova as a high-energy neutrino source (R. Abbasi et al., 2023b). However, follow-up observations of neutrino alerts do find a few interacting SNe candidates (interacting SNe in general including Ibn/IIn) (M. G. Aartsen et al., 2015b; R. Stein et al., 2025). Recently, R. Stein et al. (2025) found an interacting supernova, SN 2023uqf, in the optical follow-up observations of IceCube alert events with the Zwicky Transient Facility (ZTF), which coincided in time with the high-energy neutrino IC231004A, providing an observational evidence that interacting supernovae can serve as hadronic cosmic-ray accelerators.
@@
In view of the above arguments, a systematic search for associations between SNe IIn and neutrinos is necessary. This will help determine whether SNe IIn are effective high-energy neutrino emitters and quantify their contribution to the IceCube diffuse neutrino flux. In this work, we conduct a systematic search for spatial-temporal coincidences between SNe IIn and IceCube alert events. We use the SNe IIn classified by the ZTF - Bright Transient Survey (BTS) (D. A. Perley et al., 2020) as our supernova sample. This is a public catalog with a strict pipeline to conduct a classification for transient sources, and it conducts a magnitude-limited ($m<19\,{\rm mag}$ in either the $g$ or $r$ filter) survey for extragalactic transients in the ZTF public stream (D. A. Perley et al., 2020). We find two spatial-temporal coincidence events, namely IC231027A-SN2023syz ($z$ = 0.037) and IC250421A-SN2025cbj ($z$ = 0.0675). We calculate the model-predicted number of neutrino events for these two sources to check whether they are compatible with observations.
@@ 2.1 The Neutrino Sample
We note that R. Abbasi et al. (2023c) has provided alert events from before 2019. However, considering the start time of the ZTF operation, we focus primarily on neutrinos released after 2019, since this period better matches the survey time of the ZTF-BTS catalog. This allows alert events to occur after the SNe listed in the ZTF-BTS catalog. Therefore, for consistency, in this analysis we use only alerts listed in the GCN after 2019. We include all bronze and gold alert events in our sample, which have at least a 30% probability of astrophysical origin. In addition, in the association analysis, a good spatial localization of neutrino events is essential, as large localization errors can lead to a high rate of chance coincidences. Many previous studies (e.g., A. Plavin et al. (2020); T. Hovatta et al. (2021)) excluded alert events when the area of their 90% confidence level (C.L.) error region (denoted as $\Omega_{90}$) exceeds a certain threshold. Accordingly, in this analysis, we retain only alert events with $\Omega_{90}\leq 30\,\text{deg}^{2}$ to remove those alerts with large error regions. Our final alert sample contains a total of 138 alert events from June 19, 2019 to June 30, 2025, as shown in Fig. 1.
@@ 2.3 Crossmatch and Chance Probability
We perform a systematic search for SNe IIn coincident with neutrino events based on the following criteria: (1) the position of the SN IIn lies within the $r_{90}$ error radius of the neutrino event; (2) the arrival time of the neutrino alert event is within 100 days of the discovery date of the SN IIn. We find that SN2025cbj and SN2023syz are spatially and temporally associated with the neutrino events IC250421A and IC231027A, respectively (see Table 1 for their information). IC250421A and IC231027A arrived 61 and 38 days after the discovery date of SN 2025cbj and SN 2023syz (2025 February 19 and 2023 September 19, MJD == 60725 and 60206), respectively. The angular distances between the SNe and the central positions of their corresponding alert events are 1.79 for SN 2025cbj and 2.10 for SN 2023syz. We note that the spatial coincidence between IC250421A and SN2025cbj has already been reported in S. Garrappa et al. (2025). For IC231027A, no reports of an SN association is found; but two Fermi-LAT sources were identified within a small angular distance (< 0.3°) of the best-fit position of IC231027A (S. Buson et al., 2023). We also perform a Fermi-LAT data analysis of these two sources, as described in Appendix B. In addition, L. Pfeiffer et al. (2025) found no Fermi-LAT detection at both the position of SN2025cbj and the best-fit location of IC250421A and provided upper limits on the gamma-ray emission above 100 MeV from these two positions.
@@
$t_{\rm rise,obs}^{a}$ $M_{\rm abs}^{b}$ Redshift RA DEC Association $E_{\nu}$ Time delay
[days] [deg] [deg] [TeV] [days]
SN2023syz 10 17.58-17.58 0.037 268.85 45.22 IC231027A 191.5 38
SN2025cbj 50 19.15-19.15 0.0675 239.92 27.11 IC250421A 151.4 61
  • a The time interval between the first detection time and the peak time of the SNe.
  • b The absolute magnitude for these two sources, which can serve as indicators of the source’s emission power.
We then estimate the chance probability of obtaining two coincidence events using Monte Carlo (MC) simulations. Since IceCube is located at the South Pole and its detection sensitivity depends only on the zenith angle (M. G. Aartsen et al., 2017b, 2020), we generate mock neutrino samples by randomizing the right ascension (RA) while keeping the declination (DEC) fixed at their observed values. The arrival times of the mock neutrinos are uniformly drawn between the earliest and latest arrival times in the dataset. In addition, we generate simulated SN IIn samples by randomizing their positions within a radius of $10^{\circ}$ around their original locations [^4] The $10^{\circ}$ choice of the SNe IIn simulation follows previous works (e.g., S. Buson et al., 2022; A. V. Plavin et al., 2023; S. Buson et al., 2023; M.-X. Lu et al., 2025a), which ensure sufficient randomization while preserving the actual spatial distribution. We have also tested additional radii ranging from 4 to 20, finding that this parameter only has a small impact on the results., while requiring that each simulated source remain within the sky coverage of the ZTF-BTS field. We also apply a Kolmogorov–Smirnov (KS) test to the distributions of $\rm RA$ and $\rm DEC$ for the simulated sources to ensure that the mock samples preserve the distribution characteristics of the real sources. The mock source catalogs generated in this way has almost identical spatial distribution as the real catalog (see Apendix A).
By performing $10^{4}$ Monte Carlo simulations, we estimate the chance probability (i.e., the $p$-value for rejecting the null hypothesis) of finding at least two matches with a time delay of less than 61 days. The probability was computed using p=(M+1)/(N+1)p={(M+1)}/{(N+1)} where $M$ is the number of simulations in which the statistic (i.e., the number of matches) is greater than or equal to the observed value, and $N=10^{4}$ is the total number of MC simulations (A. C. Davison & D. V. Hinkley, 1997). We obtain a chance probability of $p\sim 0.67\%$. This result reflects that the probability of occurring two (or more) coincidence in our sample by chance is low. It is likely that the SNe and neutrino associations we have found are physically real, but the current results have not yet reached a level of statistically significant.
3 Model-predicted neutrino counts from SN 2025cbj and SN 2023syz
@@
where $R_{\rm dec}={M_{\rm ej}v_{w}}/{\dot{M}}$ is the deceleration radius, and $s=10$ is the outer slope of the ejecta density profile. Consequently, the shock velocity is $v_{\rm{sh}}(t)=dR_{\rm sh}/dt\propto t^{(-1)/(s-2)}$ for $R\leq R_{\rm dec}$ and $v_{\rm{sh}}(t)\propto t^{-1/3}$ for $R>R_{\rm dec}$.
@@
where $\alpha$ is the radial dependence index of the shock velocity, $\alpha=-1/7$ (1/2-1/2) for $R<R_{\rm dec}$ ($R>R_{\rm dec}$). $\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$ are the minimum Lorentz factor of protons, the proportion of kinetic energy converted into proton acceleration, and the proton spectral index, respectively. We follow the previous works of, e.g., T. Pitik et al. (2022); K. Murase et al. (2011), and set them as ($\gamma_{p,\rm{min}}$, $\epsilon_{p}$, $k$) = ($1,0.1,2$). The maximum Lorentz factor of protons $\gamma_{p,\rm max}$ can be calculated by requiring that the acceleration timescale is shorter than the cooling timescale. For proton cooling, we consider the cooling due to inelastic collisions and adiabatic expansion of the shocked shell. More details on the calculation of $\gamma_{p,\rm max}$ can be seen in T. Pitik et al. (2022).
@@
where $A_{\rm eff}(E_{\nu},\delta)$ is the detector effective area for which we use the effective area given in R. Abbasi et al. (2023c). For a given DEC bin, the dependence of $A_{\rm eff}$ on neutrino energy is shown in Figure 3 of R. Abbasi et al. (2023c). We determine $E_{\nu,\rm min}$ based on the lower boundary of the effective area curve. For SN2023syz (DEC = 45.22) and SN2025cbj (DEC = 27.11), the corresponding values of $E_{\nu,\rm min}$ are 1.0 TeV and 1.4 TeV, respectively.
$E_{\nu,\rm max}$ should have the same evolutionary profile as that of $E_{\rm p,max}$, and it has lower energy than $E_{\rm{p,max}}$. The value of $E_{\rm p,max}$ varies with the shock radius/time. Before and after $R_{\rm dec}$, it follows $E_{\rm p,max}\propto R^{4/7}\propto t^{1/2}$ and $E_{\rm p,max}\propto R^{-1/2}\propto t^{-1/3}$, respectively (T. Pitik et al., 2022). This indicates that $E_{\rm p,max}$ and $E_{\nu,\rm max}$ peak at $R_{\rm dec}$. For a given set of SN model parameters (i.e. $M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$, and $E_{k}$), the evolution of $E_{\nu,\rm max}$ is determined by the relative ordering of $R_{\rm dec}$, $R_{\rm bo}$, and $R_{\rm CSM}$ [^5] The parameter space for the SN model should be constrained to cases satisfying $R_{\rm bo}<R_{\rm CSM}$, since efficient particle acceleration occurs only after $R_{\rm bo}$.. If $R_{\rm bo}<R_{\rm dec}<R_{\rm CSM}$, $E_{\nu,\rm max}$ initially increases until the shock reaches $R_{\rm dec}$, after which it decreases. In the cases $R_{\rm dec}<R_{\rm bo}<R_{\rm CSM}$ and $R_{\rm bo}<R_{\rm CSM}<R_{\rm dec}$, $E_{\nu,\rm max}$ would continuously decrease and increase, respectively, throughout the shock evolution ranging from $R_{\rm bo}$ to $R_{\rm CSM}$.
To cover the broad parameter space of SNe IIn (T. Pitik et al., 2022, 2023; C. L. Ransome & V. A. Villar, 2024), we explore a wide range of values for ($M_{\rm ej}$, $M_{\rm CSM}$, $R_{\rm CSM}$). Based on ZTF-BTS observations, the adopted $E_{k}$ ranges are $(1-8)\times$ 1050 erg for SN2023syz and $(0.3-3)\times$ 1052 erg for SN2025cbj. It should be noted that, for SN2025cbj, the adopted kinetic energy values may be relatively optimistic, although it seems they can still align with the limited existing observations (see Appendix C for details).
@@
In Figure 2, we present the neutrino event rate of SN2023syz and SN2025cbj for chosen sets of model parameters. By integrating the temporal interval from the breakout time $t_{\rm{bo}}$ to the arrival dates of alert events $t_{f}$, i.e. $t_{\rm bo}=60206$ (MJD), $t_{f}=60244$ (MJD) for SN2023syz and $t_{\rm bo}=60725$ (MJD), $t_{f}=60786$ (MJD) for SN2025cbj, the total number of neutrinos can be calculated by $N_{\nu_{\mu}+\bar{\nu}_{\mu}}=\int_{t_{\rm{bo}}}^{t_{f}}dt\ \dot{N}_{\nu_{\mu}+\bar{\nu}_{\mu}}(t)$.
(a) 1-100 TeV for SN2023syz
(b) 0.1-10 PeV for SN2023syz
(c) 1-100 TeV for SN2025cbj
(d) 0.1-10 PeV for SN2025cbj
@@
Figure 2: Neutrino event rates integrated over different energy ranges as a function of the time after the shock breakout for SN2023syz and SN2025cbj, respectively. The results assume the following parameters sets: $(E_{k},R_{\rm CSM})=(8\times 10^{50}\rm erg,4\times 10^{16}cm)$ for SN2023syz and $(E_{k},R_{\rm CSM})=(3\times 10^{52}\rm erg,4\times 10^{16}cm)$ for SN2025cbj, respectively. The blue band denotes the variation arising from varying the $M_{\rm ej}$ and $M_{\rm CSM}$ parameters in the range of ($M_{\rm ej}$, $M_{\rm CSM}$) $\in$ ($1-20]\,M_{\odot}$, $1-30]\,M_{\odot}$) for SN2023syz and ($M_{\rm ej}$, $M_{\rm CSM}$) $\in$ ($_{\odot}, [170]M[1-70]\,M_{\odot}$) for SN2025cbj. When deriving the blue band, we have excluded the parameters not satisfying $t_{\rm rise}\leq t_{\rm rise,obs}\leq 1.5t_{\rm rise}$ and $t_{\rm CSM}-t_{\rm bo}$ greater than the observed time delay, please see the main text and Figure 3 for details.
@@
Figure 3 presents the model-predicted total number of muon (anti)neutrino events detected by IceCube from SN2023syz and SN2025cbj. In the figure, we have excluded those parameters (shaded region) in the parameter space according to the $t_{\rm{rise}}$ (the time interval between the first detection time [^6] Note that using the first detection time is a conservative choice, as emission is usually present before it. and the peak time of the SNe) and $t_{\rm{CSM}}-t_{\rm{bo}}$ (the time delay between $t_{\rm{bo}}$ and $t_{\rm{CSM}}$), where $t_{\rm{CSM}}$ is the time shock reaching the radius $R_{\rm CSM}$. According to the light curves of SN2023syz [^7] https://alerce.online/object/ZTF23abfglcy and SN2025cbj [^8] https://alerce.online/object/ZTF25aagbrpb, we adopt the $t_{\rm{rise,obs}}$ of 10 days and 50 days for SN2023syz and SN2025cbj, respectively. The $t_{\rm{rise}}$ can be linked to the other quantities through the following relation (S. Ginzburg & S. Balberg, 2012; T. Pitik et al., 2023),